# Exponential actions defined by vector configurations, Gale duality, and moment-angle manifolds

### v1

@@ 14.2. The nef and ample cone of a toric variety
A toric variety $X_{\Sigma}$ is defined by a lattice $N\cong\mathbb{Z}^{n}$ and a fan $\Sigma$ in $N\otimes_{\mathbb{Z}}\mathbb{R}=W^{*}$ which is rational with respect to $N$. Let $\{\mathrm{A},\mathcal{C}\}$ be the fan data of $\Sigma$, where $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ is the set of primitive generators of one-dimensional cones of $\Sigma$ and $\mathcal{C}$ is an $\mathrm{A}$-closed collection of subsets of $[m]$, see Section 4. We continue assuming that $\mathrm{A}$ spans $W^{*}$. The quotient construction [15, §5.1] identifies $X_{\Sigma}$ with the good categorical quotient $U_{\mathbb{C}}(\mathcal{K})/\!\!/G$, where $\mathcal{K}$ is the smallest simplicial complex containing $\mathcal{C}$ and $G$ is given by (14.3). The categorical quotient $U_{\mathbb{C}}(\mathcal{K})/\!\!/G$ is geometric if and only if the fan $\Sigma$ is simplicial (so that $\mathcal{C}=\mathcal{K}$).
@@
The Picard group $\mathop{\mathrm{Pic}}(X_{\Sigma})\subset\mathop{\mathrm{Cl}}(X_{\Sigma})$ consists of classes of Cartier divisors. It has a finite index in $\mathop{\mathrm{Cl}}(X_{\Sigma})$ if and only if the fan $\Sigma$ is simplicial [15, Proposition 4.2.7]. When $\Sigma$ simplicial, we have $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}=\mathop{\mathrm{missing}}{Cl}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}=V^{*}$, while in general $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}$ is a proper subspace in $V^{*}$.
@@
A torus-invariant Cartier divisor $D$ on $X_{\Sigma}$ is basepoint free if and only if the corresponding function $\varphi_{D}\colon|\Sigma|\to\mathbb{R}$ is convex [15, Theorem 6.1.7]. When the fan $\Sigma$ is complete, a torus-invariant Cartier divisor $D$ on $X_{\Sigma}$ is ample if and only if the corresponding function $\varphi_{D}\colon|\Sigma|\to\mathbb{R}$ is strictly convex [15, Theorem 6.1.14]. In particular, a toric variety $X_{\Sigma}$ is projective if and only if $\Sigma$ is the normal fan of a convex polytope $P$.
@@
The data $\{{\textbf{{w}}}_{I}\colon I\in\mathcal{C}\}$ defines a piecewise linear function $\varphi\colon|\Sigma|\to\mathbb{R}$ given by $\varphi({\textbf{{a}}})=\langle{\textbf{{a}}},{\textbf{{w}}}_{I}\rangle$ for ${\textbf{{a}}}\in\mathop{\mathrm{cone}}\mathrm{A}_{I}$. From (14.5) we obtain $\varphi({\textbf{{a}}}_{i})=-b_{i}$ for $i\in I$. Since each ${\textbf{{a}}}_{i}$ belongs to a cone, we obtain $\varphi({\textbf{{a}}}_{i})=-b_{i}$ for any $i$. Identity (14.5) also implies that
@@
so that $\varphi$ is convex. Since $\varphi({\textbf{{a}}}_{i})$ is rational for any $i$, multiplying by a common denominator we achive that $\varphi$ is integer on $N$. It follows that $\gamma=\sum_{i=1}^{m}b_{i}[D_{i}]=-\sum_{i=1}^{m}\varphi({\textbf{{a}}}_{i})[D_{i}]$ is a basepoint free divisor class up to a positive multiple.
@@
When $X_{\Sigma}$ is not simplicial, $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ cannot have full dimension in $V^{*}$, as $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})\subset\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}$ and $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}$ is a proper subspace in $V^{*}$. However, $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ can have full dimension in $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}$. When $X_{\Sigma}$ is projective, $\mathop{\mathrm{missing}}{Ample}(X_{\Sigma})$ is the relative interior of $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ (or the interior of $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ in $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}$). On the other hand, $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ may have full dimension in $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}$ with $X_{\Sigma}$ still not being projective. To illustrate this, we elaborate on an example due to Fujino [17], see also [15, Example 6.3.26].
@@
Figure 4.
It can be shown that $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})=\mathbb{Z}\langle 3(\gamma_{1}+\gamma_{4})\rangle$, so $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ has full dimension in $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})$, yet $3(\gamma_{1}+\gamma_{4})$ is a Cartier divisor class in the interior of $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ which is not ample.
@@ 14.3. The Kähler cone and symplectic reduction
When $X_{\Sigma}$ is complete and smooth, the interior of $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ consists of Kähler classes in $V^{*}=H^{2}(X_{\Sigma};\mathbb{R})$ and is known as the Kähler cone. It is empty when $X_{\Sigma}$ is nonprojective. The space $V^{*}$ contains the lattice $L^{*}=\mathbb{Z}\langle\gamma_{1},\ldots,\gamma_{m}\rangle=\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})=H^{2}(X_{\Sigma};\mathbb{Z})$, and the lattice points in the interior of $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ correspond to Kähler classes coming from ample divisors on $X_{\Sigma}$.
A Hamiltionian toric manifold $M$ can be reconstructed from its moment polytope $P$ using symplectic reduction. We outline the construction below; the details can be found in [2, Chapter 6] and [20, Chapter 1].
Construction 14.3 (Hamiltonian toric manifolds as symplectic quotients).
@@
where $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ is the Gale dual configuration of $\Gamma$ and $\delta=\sum_{i=1}^{m}b_{i}\gamma_{i}$, see Theorem 8.4. (Note that $P$ is defined by $(\Gamma,\delta)$ up to a translation in $W$.) The normal fan $\Sigma_{P}$ of $P$ is defined by the data $(\mathcal{K}_{P},\mathrm{A})$, and it is nonsingular if and only if the action of $K$ on $\mu_{\Gamma}^{-1}(\delta)$ is free, see Proposition 11.4. A polytope $P$ with a nonsingular normal fan is called Delzant. (Note that the vertices of a Delzant polytope do not necessarily lie in the lattice $M=N^{*}\subset W$.)
@@
To prove the first statement we need to check that each $G$-orbit of $U_{\mathbb{C}}(\mathcal{K}_{P})$ intersects $\mu_{\Gamma}^{-1}(\delta)$ at a single $K$-orbit. Using the product decomposition $G=\mathbb{C}^{\times}_{L}=V\times T_{L}=V\times K$, we need to the check that each $V$-orbit of $U(\mathcal{K}_{P})$ intersects $\mu_{\Gamma}^{-1}(\delta)$ at a single point. This is proved in Theorem 8.4 (4). The second theorem follows from the Delzant theorem, as a $T_{N}$-equivariant symplectic structure on $X_{\Sigma_{P}}$ is determined by its moment polytope $P$ uniquely. ∎
There is a generalisation of Theorem 14.4 to nonrational simple polytopes $P$ [5, Theorem 3.2], which relates the algebraic and symplectic quasifolds associated with $P$ and a quasilattice containing the normal configuration $\mathrm{A}$.

### v2

@@ 14.2. The nef and ample cone of a toric variety
A toric variety $X_{\Sigma}$ is defined by a lattice $N\cong\mathbb{Z}^{n}$ and a fan $\Sigma$ in $N\otimes_{\mathbb{Z}}\mathbb{R}=W^{*}$ which is rational with respect to $N$. Let $\{\mathcal{C},\mathrm{A}\}$ be the fan data of $\Sigma$, where $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ is the set of primitive generators of one-dimensional cones of $\Sigma$ and $\mathcal{C}$ is an $\mathrm{A}$-closed collection of subsets of $[m]$, see Section 4. We continue assuming that $\mathrm{A}$ spans $W^{*}$. The quotient construction [16, §5.1] identifies $X_{\Sigma}$ with the good categorical quotient $U_{\mathbb{C}}(\mathcal{K})/\!\!/G$, where $\mathcal{K}$ is the smallest simplicial complex containing $\mathcal{C}$ and $G$ is given by (14.3). The categorical quotient $U_{\mathbb{C}}(\mathcal{K})/\!\!/G$ is geometric if and only if the fan $\Sigma$ is simplicial (so that $\mathcal{C}=\mathcal{K}$).
@@
The Picard group $\mathop{\mathrm{Pic}}(X_{\Sigma})\subset\mathop{\mathrm{Cl}}(X_{\Sigma})$ consists of classes of Cartier divisors. It has a finite index in $\mathop{\mathrm{Cl}}(X_{\Sigma})$ if and only if the fan $\Sigma$ is simplicial [16, Proposition 4.2.7]. When $\Sigma$ simplicial, we have $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}=\mathop{\mathrm{missing}}{Cl}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}=V^{*}$, while in general $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}$ is a proper subspace in $V^{*}$.
@@
A torus-invariant Cartier divisor $D$ on $X_{\Sigma}$ is basepoint free if and only if the corresponding function $\varphi_{D}\colon|\Sigma|\to\mathbb{R}$ is convex [16, Theorem 6.1.7]. When the fan $\Sigma$ is complete, a torus-invariant Cartier divisor $D$ on $X_{\Sigma}$ is ample if and only if the corresponding function $\varphi_{D}\colon|\Sigma|\to\mathbb{R}$ is strictly convex [16, Theorem 6.1.14]. In particular, a toric variety $X_{\Sigma}$ is projective if and only if $\Sigma$ is the normal fan of a convex polytope $P$.
@@
The data $\{{\textbf{{w}}}_{I}\colon I\in\mathcal{C}\}$ define a piecewise linear function $\varphi\colon|\Sigma|\to\mathbb{R}$ given by $\varphi({\textbf{{a}}})=\langle{\textbf{{a}}},{\textbf{{w}}}_{I}\rangle$ for ${\textbf{{a}}}\in\mathop{\mathrm{cone}}\mathrm{A}_{I}$. From (14.5) we obtain $\varphi({\textbf{{a}}}_{i})=-b_{i}$ for $i\in I$. Since each ${\textbf{{a}}}_{i}$ belongs to a cone, we obtain $\varphi({\textbf{{a}}}_{i})=-b_{i}$ for any $i$. Identity (14.5) also implies that
@@
so that $\varphi$ is convex. Since $\varphi({\textbf{{a}}}_{i})$ is rational for any $i$, multiplying by a common denominator we achieve that $\varphi$ is integer on $N$. It follows that $\gamma=\sum_{i=1}^{m}b_{i}[D_{i}]=-\sum_{i=1}^{m}\varphi({\textbf{{a}}}_{i})[D_{i}]$ is a basepoint free divisor class up to a positive multiple.
@@
When $X_{\Sigma}$ is not simplicial, $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ cannot have full dimension in $V^{*}$, as $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})\subset\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}$ and $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}$ is a proper subspace in $V^{*}$. However, $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ can have full dimension in $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}$. When $X_{\Sigma}$ is projective, $\mathop{\mathrm{missing}}{Ample}(X_{\Sigma})$ is the relative interior of $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ (or the interior of $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ in $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}$). On the other hand, $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ may have full dimension in $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}$ with $X_{\Sigma}$ still not being projective. To illustrate this, we elaborate on an example due to Fujino [20], see also [16, Example 6.3.26].
@@
Figure 4. A non-projective complete toric variety with the nef cone of full dimension in the Picard group
It can be shown that $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})=\mathbb{Z}\langle 3(\gamma_{1}+\gamma_{4})\rangle$, so $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ has full dimension in $\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})\otimes_{\mathbb{Z}}\mathbb{R}$, yet $3(\gamma_{1}+\gamma_{4})$ is a Cartier divisor class in the interior of $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ which is not ample.
@@ 14.3. The Kähler cone and symplectic reduction
When $X_{\Sigma}$ is complete and smooth, the interior of $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ consists of Kähler classes in $V^{*}=H^{2}(X_{\Sigma};\mathbb{R})$ and is known as the Kähler cone. It is empty when $X_{\Sigma}$ is nonprojective. The space $V^{*}$ contains the lattice $L^{*}=\mathbb{Z}\langle\gamma_{1},\ldots,\gamma_{m}\rangle=\mathop{\mathrm{missing}}{Pic}(X_{\Sigma})=H^{2}(X_{\Sigma};\mathbb{Z})$, and lattice points in the interior of $\mathop{\mathrm{missing}}{Nef}(X_{\Sigma})$ correspond to Kähler classes coming from ample divisors on $X_{\Sigma}$.
A Hamiltonian toric manifold $M$ can be reconstructed from its moment polytope $P$ using symplectic reduction. We outline the construction below; the details can be found in [2, Chapter 6] and [23, Chapter 1].
Construction 14.4 (Hamiltonian toric manifolds as symplectic quotients).
@@
where $\mathrm{A}=\{{\textbf{{a}}}_{1},\ldots,{\textbf{{a}}}_{m}\}$ is the Gale dual configuration of $\Gamma$ and $\delta=\sum_{i=1}^{m}b_{i}\gamma_{i}$, see Theorem 8.4. (Note that $P$ is defined by $(\Gamma,\delta)$ up to a translation in $W$.) The normal fan $\Sigma_{P}$ of $P$ is defined by the triangulated configuration $(\mathcal{K}_{P},\mathrm{A})$, and it is nonsingular if and only if the action of $K$ on $\mu_{\Gamma}^{-1}(\delta)$ is free, see Proposition 11.4. A polytope $P$ with a nonsingular normal fan is called Delzant. (Note that the vertices of a Delzant polytope do not necessarily lie in the lattice $M=N^{*}\subset W$.)
@@
To prove the first statement we need to check that each $G$-orbit of $U_{\mathbb{C}}(\mathcal{K}_{P})$ intersects $\mu_{\Gamma}^{-1}(\delta)$ at a single $K$-orbit. Using the product decomposition $G=\mathbb{C}^{\times}_{L}=V\times T_{L}=V\times K$, we need to the check that each $V$-orbit of $U(\mathcal{K}_{P})$ intersects $\mu_{\Gamma}^{-1}(\delta)$ at a single point. This is proved in Theorem 8.4 (4). The second statement follows from the Delzant theorem, as a $T_{N}$-equivariant symplectic structure on $X_{\Sigma_{P}}$ is determined by its moment polytope $P$ uniquely. ∎
The complex structure on the algebraic quotient $U_{\mathbb{C}}(\mathcal{K}_{P})/G$ is compatible with the symplectic structure on the symplectic quotient $\mu_{\Gamma}^{-1}(\delta)/K$, so that the symplectic form $\omega^{\prime}$ on $\mu_{\Gamma}^{-1}(\delta)/K$ is Kähler. See [23, Appendix 1, Theorem 1.5].
There is a generalisation of Theorem 14.5 to irrational simple polytopes $P$ [5, Theorem 3.2], which relates the complex and symplectic quasifolds associated with $P$ and a quasilattice containing the normal configuration $\mathrm{A}$.